# Enhancing Robustness to Noise Corruption for Point Cloud Model via Spatial Sorting and Set-Mixing Aggregation Module

### v1

@@ 3.1 Point-Based Network Analysis
In this paper, we propose an MLP-based order-sensitive module, amalgamating the functions of both spatial relation learning and shape feature aggregation, named Set-Mixer. As illustrated in Fig. 1(d), the challenge of maintaining permutation invariance is addressed through spatial-aware point sorting, and an MLP is employed for feature aggregation while simultaneously constructing relative spatial information among all points. Through spatial-aware sorting, we integrate spatial information into the order of the entire point set, rather than embedding point-wise position information in traditional methods. Therefore, the proposed module directly learns relations among all points and employs non-symmetric aggregation, which effectively diminishes the impact of a minor number of noisy points.
@@ 3.3 Spatial Aware Sorting
$\mathbf{Angle}_{(x,v_{ref})}=(x\cdot\mathbf{v}_{ref})/\|x\|),\vspace{-0.5em}$ (10)
@@ 3.4 Set-Mixer Network
It is noteworthy that we adjust the operation of grouping local point sets. The original version of PointNet++ incorporates a process of generating local reference frames by subtracting the position of the query point $cp$ from each point, expressed as:
$\mathbf{P}_{local}=\mathbf{P}-cp.$ (14)
It then hierarchically embeds the relative coordinates of higher levels or global coordinates of query points into the network. This operation introduces the possibility of recurrently feeding noise information into the network, potentially directly influencing high-level features and thus increasing the difficulty of the network in mitigating the impact of outliers.
@@
The MLP layer $m$ consists of three Fully Connected (FC) layers, facilitating channel-wise communication among all points. It aggregates the input feature matrix $\mathbf{f}_{\mathcal{O}}\in\mathbb{R}^{k\times 3c}$ into $\mathbf{f}_{m}\in\mathbb{R}^{d\times 3c}$, where $d$ represents the channel dimension of $m$. Given that the sorting process was implemented multiple times, effectively increasing the feature matrix size, an additional FC layer $r$ was employed to reduce the feature channels from $d*3*c$ to $2c$ to maintain the output feature size. Additionally, $m$ incorporates two dropout layers, which force the network to avoid over-reliance on specific points, such as those with extreme coordinate values. The mixing module comprehensively considers the relations among all points during aggregation, maximizing the mitigation of the impact of noisy points.
@@ 4.1 Model Setting
The proposed Set-Mixer module is incorporated into PointNet++ [23] structure and the detailed structure can be found in the appendix. Our Set-Mixer is capable of conducting multiple sorting operations. To maximize the acquisition of spatial information, we utilize the three orthogonal directions x, y, and z-axes for APS sorting. Similarly, for PCS sorting, the XY-plane, YZ-plane, and XZ-plane are employed, with the x-axis, y-axis, and z-axis serving as the corresponding starting reference vector $v_{ref}$, respectively. EDS, on the other hand, is limited to a single iteration. However, despite this constraint, we included tests in our ablation study to assess its combined impact with other sorting methods.
@@ 4.2 Dataset and Evaluation Criteria
Following the ModelNet40-C benchmark, our model was trained on the clean ModelNet40 dataset and employed the corrupted ModelNet40-C as the test set. ModelNet40 [35] is a synthetic dataset widely utilized in the field of point cloud analysis, which has been sampled from 40 distinct categories of CAD models. In the official rendition, the dataset is partitioned into 9,843 samples for training and 2,468 samples for testing. ModelNet40-C [28] is derived by introducing corruption to the test set of ModelNet40. It defines 15 distinct corruption types, with each category consisting of five variations in Density, Noise, and Transformation. Furthermore, the dataset is generated with five severity levels for each corruption type. We selected Uniform, Gaussian, Impulse, Upsampling and Background from the Noise category for evaluation. All training and testing experiments are done on a GeForce RTX 3090 GPU.
The error rate (ER), class-wise mean error rate (mER) and Relative mean CE (RmCE), are employed as the metrics. Following the definition of ModelNet40-C, $\mathbf{ER}^{f}_{s,cr}$ denotes the error rate of classifier $f$ on corruption $cr$ with severity $s$, $\mathbf{ER}^{f}_{s}$ and $\mathbf{ER}^{f}_{\text{noise}}$ denote the mean error rate of severity $s$ and the mean error rate of 5 noise corruption types, respectively.
RmCE is introduced by PointCloud-C, which is designed for measuring the performance drop compared to the clean test set. PointNet++ with the smallest performance drop among typical models are selected as the baseline model (bm) and the metric can be formulated as:
$\mathbf{RmCE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{ER}^{f}_{\text{noise}}-\mathbf{ER}^{f}_{\text{clean}}}{\mathbf{ER}^{bm}_{\text{noise}}-\mathbf{ER}^{bm}_{\text{clean}}}.$ (17)
@@
4.3 Performance Analysis
The comparison of our proposed Set-Mixer and other representative models are illustrated in Tab. 1, demonstrating the superior robustness of the Set-Mixer to noisy data. The proposed method exhibits the lowest error rate across all categories, particularly excelling in the Impulse and Background types, surpassing the PointNet++ by $25.4\%$ and $8.2\%$, respectively. This is attributed to the fact that these two types of corruption introduce noisy points with extreme values. As previously discussed, such values will be captured by max-pooling methods, substantially contaminating the aggregated feature. In contrast, their impact on the Set-Mixer structure is minimal. Furthermore, it is worth mentioning that although the proposed model exhibits a relatively higher error rate on the clean dataset, the disparity introduced by noisy corruption is significantly smaller compared to other models. The performance on the clean dataset could potentially be further refined through the implementation of more sophisticated sorting strategies for spatial information generation.
@@ 4.4 Data Augmentation Strategies
The proposed Set-Mixer enhances robustness by optimizing the aggregation module, thereby allowing it to effectively integrate with well-established generic techniques to further improve its robustness. By employing augmentation strategies, at the expense of increased training time, performance can be further enhanced. Given the remarkable efficacy of PointCutMix-R, PointCutMix-K [42], and RSMix [15] in enhancing the robustness of conventional models, we assessed their impact when used with Set-Mixer. As shown in Tab. 2, augmentation methods efficiently reduce the error rate of all methods, with Set-Mixer achieving the lowest error rate. It is noteworthy that EPiC uses WolfMix [24], which combines PointWolf [14] with RSMix, but this additional operation does not significantly improve robustness to noise. The error rate under severity 5 is also calculated and it shows the noise robustness advantage of Set-Mixer becomes even more pronounced as the noise level increases.
Table 3: Classification error rate on ModelNet40-C with different sorting strategies.
@@
Uniform Gaus. Impulse Upsamp. Bg.
dp 0.0 8.1 10.3 0.15 10.6 10.7 9.8 10.0 10.3
dp 0.1 8.7 10.5 0.12 10.5 10.7 10.0 10.5 10.6
dp 0.2 8.1 10.0 0.13 10.5 10.1 9.7 9.4 10.4
dp 0.3 8.5 10.2 0.12 10.1 10.0 9.5 10.3 11.1
dp 0.0 8.8 10.9 0.11 10.9 10.9 10.2 10.8 11.5
dp 0.2 8.8 10.5 0.12 10.7 10.6 9.9 10.1 11.2
4.5 Sorting Strategy
In the proposed Set-Mixer, we perform feature sorting three times independently along the xyz-axis to enrich spatial information. We also evaluated alternative sorting strategies, including sorting once or twice, with the comprehensive results illustrated in Tab. 3.
As demonstrated by the table, all sorting methods exhibit superior performance compared to PointNet++. Notably, conducting multiple sorting operations yields a significant improvement. In general, increasing the number of sorting operations would enrich the spatial information of point sets, leading to enhanced performance and robustness.
4.6 Dropout and Layer Normalization
Two dropout layers and a layer normalization layer are employed in the Set-Mixer, preventing the layer over-reliance on specific points and enhancing training stability. To further analyze the impact of the those two operations, we conducted an ablation experiment. The results, presented in Tab. 4, indicate that the dropout layer exerts minimal influence on performance when assessed on a clean test set but exhibits a more pronounced impact on noisy test sets. Remarkably, the proposed model demonstrates substantial robustness to noise even in the absence of the dropout layer. Besides, we observe that layer normalization effectively enhances accuracy on both clean and corrupted datasets. Nonetheless, our model still outperforms other models without layer normalization, indicating that the primary contributing factor is the Set-Mixer’s architecture, which is designed to effectively extract geometric shape information.
Query Point Spatial Center Max Pooling Mean Pooling Mixer w/o sorting
$\text{ER}_{\text{clean}}$ 8.5 8.1 8.5 12.8 10.6
$\text{ER}_{\text{noise}}$ 10.8 10.0 15.7 18.9 13.4
RmCE 0.16 0.13 0.50 0.42 0.19
Figure 5: Visualization of query points and spatial center points of layer 1 ((((512 sets, 32 neighbours)))) and layer 2 ((((128 sets, 64 neighbours)))) of Set-Mixer. Query points are depicted in blue, while the spatial centers are marked in orange.
4.7 Analysis of Spatial Center and Sorting Strategy Analysis
We conducted additional comparative experiments using Set-Mixer-APS to substantiate the efficacy of sorting based on the spatial center. As shown in Tab. 5, leveraging the spatial center outperforms the query point center. Remarkably, even without incorporating the spatial center, our model’s sorting strategy substantially enhances robustness. In Fig. 5, we visualize both the query points and the spatial centers generated by Set-Mixer in layers 1 and 2. Through this visualization, we can observe that the spatial center not only provides a more effective representation of the spatial positions of point sets but also exhibits greater stability.
Furthermore, we tested the model using a pooling function and the mixer without the sorting operation. The results indicate that the mixer structure is more robust, while the sorting operation effectively improves performance.
5 Conclusion
Appendix
Appendix 0.A Model Structure
[…]
Appendix 0.B Noisy Feature Visualization

### v2

# Enhancing Robustness to Noise Corruption for Point Cloud Recognition via Spatial Sorting and Set-Mixing Aggregation Module

@@ 3.1 Point-based Network Analysis
In this paper, we propose an MLP-based order-sensitive module, amalgamating the functions of both spatial relation learning and shape feature aggregation, named Set-Mixer. As illustrated in Fig. 1(d), the challenge of maintaining permutation invariance is addressed through spatial-aware point sorting, and an MLP is employed for feature aggregation while simultaneously constructing relative spatial information among all points. Through spatial-aware sorting, we integrate spatial information into the order of the entire point set, rather than embedding point-wise position information. Therefore, the proposed module directly learns relations among all points and employs non-symmetric aggregation, which effectively diminishes the impact of a minor number of noisy points.
@@ 3.3 Spatial Aware Sorting
$\mathbf{Angle}_{(x,v_{ref})}=(x\cdot\mathbf{v}_{ref})/\|x\|),$ (10)
@@ 3.4 Set-Mixer Network
It is noteworthy that we adjust the operation of grouping local point sets. The original version of PointNet++ incorporates a process of generating local reference frames by subtracting the position of the query point $cp$ from each point, expressed as: $\mathbf{P}_{local}=\mathbf{P}-cp.$ It then hierarchically embeds the relative coordinates of higher levels or global coordinates of query points into the network. This operation introduces the possibility of recurrently feeding noise information into the network, directly influencing high-level features and thus increasing the difficulty of the network in mitigating the impact of outliers.
@@
The MLP layer $m$ consists of three Fully Connected (FC) layers, facilitating channel-wise communication among all points. It aggregates the input feature matrix $\mathbf{f}_{\mathcal{O}}\in\mathbb{R}^{k\times 3c}$ into $\mathbf{f}_{m}\in\mathbb{R}^{d\times 3c}$, where $d$ represents the channel dimension of $m$. Given that the sorting process was implemented multiple times, effectively increasing the feature matrix size, an additional FC layer $r$ was employed to reduce the feature channels from $d*3*c$ to $2c$ to maintain the output feature size. Additionally, $m$ incorporates two dropout layers, which force the network to avoid over-reliance on specific points, such as those with extreme coordinate values. The mixing module comprehensively considers the relations among all points for aggregation, maximizing the mitigation of the impact of noisy points.
@@ 4.1 Model Setting
The proposed Set-Mixer module is incorporated into PointNet++ [22] structure and the Set-Mixer is capable of conducting multiple sorting operations. To maximize the acquisition of spatial information, we utilize the three orthogonal directions x, y, and z-axes for APS sorting. Similarly, for PCS sorting, the XY-plane, YZ-plane, and XZ-plane are employed, with the x-axis, y-axis, and z-axis serving as the corresponding starting reference vector $v_{ref}$, respectively.
@@ 4.2 Dataset and Evaluation Criteria
Following the ModelNet40-C [27] and PointCloud-C [23] benchmark, our model was trained on the clean ModelNet40 [35] dataset and employed the corrupted test sets for evaluation. Both of ModelNet40-C and PointCloud-C are derived by introducing corruption to the test set of ModelNet40. ModelNet40-C defines 15 distinct corruption types, with each category comprising five variations related to Density, Noise, and Transformation. Additionally, the dataset incorporates five levels of severity for each corruption type. For our evaluation, we selected Uniform, Gaussian, Impulse, Upsampling, and Background from the Noise category. For PointCloud-C, we focused on Jitter, Drop Global/Local, and Add Global/Local, excluding Scale and Rotate as they are unrelated to noise robustness. Besides, we also employ the real-world dataset ScanObjectNN [29] (PB_T50_RS variant) for evaluation and created a corrupted ScanObjectNN-C set, with the method for generating corrupted samples from ModelNet40-C [27].
The error rate (ER), class-wise mean error rate (mER) and Relative mean CE (RmCE), are employed as the metrics. Following the definition of ModelNet40-C, $\mathbf{ER}^{f}_{s,cr}$ denotes the error rate of classifier $f$ on corruption $cr$ with severity $s$, $\mathbf{ER}^{f}_{s}$ and $\mathbf{ER}^{f}_{\text{noise}}$ denote the mean error rate of severity $s$ and the mean error rate of 5 noise corruption types, respectively. RmCE is introduced by PointCloud-C, which is designed for measuring the performance drop compared to the clean test set. PointNet++ with the smallest performance drop among typical models are selected as the baseline and the metric can be formulated as:
$\mathbf{RmCE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{ER}^{f}_{\text{noise}}-\mathbf{ER}^{f}_{\text{clean}}}{\mathbf{ER}^{PointNet++}_{\text{noise}}-\mathbf{ER}^{PointNet++}_{\text{clean}}}.$ (16)
4.3 Performance Analysis
The comparison of our proposed Set-Mixer and other representative models on ModelNet40-C are illustrated in Table 1, demonstrating the superior robustness of the Set-Mixer to noisy data. The proposed method exhibits the lowest error rate across all categories, particularly excelling in the Impulse and Background types, surpassing the PointNet++ by $25.4\%$ and $8.2\%$, respectively. This is attributed to the fact that these two types of corruption introduce noisy points with extreme values. As previously discussed, such values will be captured by max-pooling methods, substantially contaminating the aggregated feature. In contrast, their impact on the Set-Mixer structure is minimal. Furthermore, it is worth mentioning that although the proposed model exhibits a relatively higher error rate on the clean dataset, the disparity introduced by noisy corruption is significantly smaller compared to other models. The performance drop on the clean dataset have also been observed in CSI [36] and CP++ [16]. In our method, this issue may stem from the difficulty of restoring positional information through the current sorting strategies, and it could potentially be further refined through the implementation of more sophisticated sorting strategies for spatial information generation.
Table 2 presents the performance of the Set-Mixer on PointCloud-C and ScanObjectNN-C, together with its model size and inference computational cost. While the Set-Mixer module increases model size, its computational cost remains low. This is because we utilize global coordinates, so that the computational load of the first set abstraction can be significantly reduced with code optimizations.
@@ 4.3 Performance Analysis
Table 2: Model efficiency and performance on PointCloud-C and ScanObjectNN-C.
Typical Model PointCloud-C ScanObjectNN-C Param. FLOPs
$\text{ER}_{\text{clean}}$ $\text{ER}_{\text{noise}}$ $\text{ER}_{\text{clean}}$ $\text{ER}_{\text{noise}}$
PointNet++ [22] 7.0 27.2 22.2 51.9 1.48M 1.68G
GDANet [40] 6.6 20.3 23.3 52.4 0.93M 9.66G
PointMLP [20] 6.5 29.7 15.6 54.0 13.24M 15.67G
EiPC-PCT [15] 6.6 16.5 - - 8.82M 10.61G
CP++-DGCNN [16] 8.6 25.2 - - 1.81M 6.88G
Set-Mixer-APS 8.1 16.4 22.8 37.4 8.02M 1.76G
Set-Mixer-PCS 9.6 16.5 22.8 37.6 8.02M 1.76G
Set-Mixer-EDS 8.2 29.5 30.7 55.7 3.50M 0.71G
@@ 4.4 Data Augmentation Strategies
The proposed Set-Mixer enhances robustness by optimizing the aggregation module, thereby allowing it to effectively integrate with well-established generic techniques to further improve its robustness. By employing augmentation strategies, at the expense of increased training time, performance can be further enhanced. Given the remarkable efficacy of PointCutMix-R, PointCutMix-K [44], and RSMix [14] in enhancing the robustness of conventional models, we assessed their impact when used with Set-Mixer. As shown in Table 3, augmentation methods efficiently reduce the error rate of all methods, with Set-Mixer achieving the lowest error rate. It is noteworthy that EPiC uses WolfMix [23], which combines PointWolf [13] with RSMix, but this additional operation does not significantly improve robustness to noise. The error rate under severity 5 is also calculated and it shows the noise robustness advantage of Set-Mixer becomes even more pronounced as the noise level increases.
Uniform Gaus. Impulse Upsamp. Bg.
dp 0.0 8.1 10.3 0.15 10.6 10.7 9.8 10.0 10.3
dp 0.1 8.7 10.5 0.12 10.5 10.7 10.0 10.5 10.6
dp 0.2 8.1 10.0 0.13 10.5 10.1 9.7 9.4 10.4
dp 0.3 8.5 10.2 0.12 10.1 10.0 9.5 10.3 11.1
dp 0.0 8.8 10.9 0.11 10.9 10.9 10.2 10.8 11.5
dp 0.2 8.8 10.5 0.12 10.7 10.6 9.9 10.1 11.2
4.5 Dropout and Layer Normalization
Two dropout layers and a layer normalization layer are employed in the Set-Mixer, preventing the layer over-reliance on specific points and enhancing training stability. To further analyze the impact of the those two operations, we conducted an ablation experiment. The results, presented in Table 4, indicate that the dropout layer exerts minimal influence on performance when assessed on a clean test set but exhibits a more pronounced impact on noisy test sets. Remarkably, the proposed model demonstrates substantial robustness to noise even in the absence of the dropout layer. Besides, we observe that layer normalization effectively enhances accuracy on both clean and corrupted datasets. Nonetheless, our model still outperforms other models without layer normalization, indicating that the primary contributing factor is the Set-Mixer’s architecture, which is designed to effectively extract geometric shape information.
Query Point Spatial Center Max Pooling Mean Pooling Mixer w/o sorting
$\text{ER}_{\text{clean}}$ 8.5 8.1 8.5 12.8 10.6
$\text{ER}_{\text{noise}}$ 10.8 10.0 15.7 18.9 13.4
RmCE 0.16 0.13 0.50 0.42 0.19
4.6 Analysis of Spatial Center and Sorting Strategy Analysis
We conducted additional comparative experiments using Set-Mixer-APS to substantiate the efficacy of sorting based on the spatial center. As shown in Table 5, leveraging the spatial center outperforms the query point center. Remarkably, even without incorporating the spatial center, our model’s sorting strategy substantially enhances robustness. Furthermore, we tested the model using a pooling function and the mixer without the sorting operation. The results indicate that the mixer structure is more robust, while the sorting operation effectively improves performance.
5 Conclusion
Appendix
Appendix 0.A Model Structure
[…]
Appendix 0.B Sorting Strategy
In the proposed Set-Mixer, we perform feature sorting three times independently along the xyz-axis to enrich spatial information. We also evaluated alternative sorting strategies, including sorting once or twice, with the comprehensive results illustrated in Table A6.
As demonstrated by Table A6, all sorting methods exhibit superior performance compared to PointNet++. Notably, conducting multiple sorting operations yields a significant improvement. In general, increasing the number of sorting operations would enrich the spatial information of point sets, leading to enhanced performance and robustness.
Table A6: Classification error rate on ModelNet40-C with different sorting strategies.
@@ Appendix 0.B Sorting Strategy
Appendix 0.C Spatial Center Visualization
In Fig. A5, we visualize both the query points and the spatial centers generated by Set-Mixer in layers 1 and 2. Through this visualization, we can observe that the spatial center not only provides a more effective representation of the spatial positions of point sets but also exhibits greater stability.
Appendix 0.D Noisy Feature Visualization
Figure A5: Visualization of query points and spatial center points of layer 1 ((((512 sets, 32 neighbours)))) and layer 2 ((((128 sets, 64 neighbours)))) of Set-Mixer. Query points are depicted in blue, while the spatial centers are marked in orange.